Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

ADDRESS HERE

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Dublin University Rifle Club Sponsorship  1st November 2001

A Chara,

The Dublin University Rifle Club is one of the oldest, largest and most successful sports clubs in Trinity College today. We celebrate the 40^{th} anniversary of the founding of the modern club this year, and last year saw the 160^{th} anniversary of the original college rifle club. We currently have 398 members, making us the largest sporting club in the College. DURC members act as part of the National Target Shooting Association committee and we have been requested by the NTSA to manage the logistics for the Air Rifle Nationals for the last two years. This past year alone, DURC members have competed in 27 national-level competitions and four international level competitions, including two World Cups; we won a total of 38 prizes at these competitions. DURC has had at least one competitor qualify for the Irish National Squad every year for the past twenty years; at present we have four, including two juniors. This year’s Munich World Cup saw a DURC member acting as Irish Team Manager. We have a long history of international competition, and an excellent reputation both here and abroad. A DURC competitor was the first Irish competitor to win a gold medal abroad and DURC competitors have represented Ireland at international level many times. We also have a healthy public image, with many articles in national, local and College newspapers and magazines, as well as occasional television coverage.

However, the grants which the College provide for the running of the club do not cover all of our expenses, and are limited in their coverage - for example, travel to competitions outside Ireland, prizes for competitions or office expenses are not covered by these grants. As such, we seek external sponsorship wherever possible. We feel that our club is a prime opportunity for corporate sponsorship given our reputation and history.

The prime events for which we are seeking sponsorship are the DURC 10m Air Rifle Open and the DURC 50m Free Rifle Open, our main external competitions during the year. Both of these events are Olympic disciplines, and usually have upwards of forty senior competitors and twenty junior competitors, from all over the country, though that number has been rising steadily in recent years. They receive coverage in national, local and college newsletters. The opens are also open to competitors from abroad, and this year we are hoping to invite competitors from the Isle of Man and from the UK. We would welcome an “¡COMPANY NAME¿ Cup” in both competitions, with the ¡COMPANY NAME¿ logo on all prizes given out.

Alternatively, sponsorship could take the form of providing sports clothing for the DURC squad with your corporate logo or funding the formal events held during the year such as the 40^{th} anniversary formal dinner function (to be held in Trinity College to mark the actual anniversary of the club’s founding).

I would be grateful for the opportunity to discuss this proposal further. I can be contacted at the above addresses; I look forward to hearing from you.

Misé le meas,

Mark Dennehy, DURC Secretary